graph f(x)=-x^2+4x-3, labeling the y-intercept, vertex, and axis of symmetry
please help.

Answer by ewatrrr(24785)


Hi,
f(x)=-x^2+4x-3
y-intercept when x = 0 Pt(0-3)
x-intercept when y = 0
y = -x^2 + 4x - 3 = 0 Completing Square to put into vertex form
y = -[x^2 - 4x] - 3 = 0
y = -[(x-2)^2 - 4] - 3 = 0
y = -(x-2)^2 + 4 - 3 = 0 Vertex Pt (2,1) axis of symmetry is x = 2